A carnot engine with a cold body temperature of 17°celcius has efficiency of 50% by how much should the temperature of its iot body be changed to increase the efficiency to 60% when will its efficiency be 100%

Grade:11

2 Answers

Arun
4 years ago
 
1case:TE=290k ,n=50%=1/2,TE=?      n=1-T2/TE=
1/2=1-290/TE
So T1=580k
Case
Now T1'=T1+418=580+418=998k
Now n'=1-T2/T1'==1-290/998=708/998
n'=708/998×100=70
Vikas TU
4 years ago